OpenStudy (anonymous):

Find the coordinates for the center for this circle. (X+3)^2 + (y-5)^2 = 25

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ObeySoph▲

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

General form of a circle:\[\Large\rm (x-\color{orangered}{h})^2 + (y-\color{royalblue}{k})^2 = r^2\]Where the center of the circle is at \(\Large\rm (\color{orangered}{h},~\color{royalblue}{k})\). We were given:\[\Large\rm (x+3)^2 + (y-5)^2 = r^2\]We need subtraction in the x so we can clearly see our h value. I'll write it as minus a negative,\[\Large\rm (x-\color{orangered}{(-3)})^2 + (y-\color{royalblue}{5})^2 = r^2\]Can you figure out the center coordinates from there?

OpenStudy (anonymous):

Alright thank you a lot, would it be (-3, 5)

zepdrix (zepdrix):

Yes, good job.

OpenStudy (anonymous):

Alright thanks for your help